A college has 25000 students. Based on student personal information data, approximately 45 percent are Democrats. A simple random sample of 900 households were selected for a survey. The percentage of Democrats in the sample will be around ___ give or take ____ or so.
Estimate the chance that between 42.45 percent and 47.55 percent of the students in the sample are Democrats. ___

Answer by Boreal(15235)

give or take means margin of error is 1/sqrt(n) or 1/30 or 0.033 or 3.3%
so it would be around 42-48 %


This is not exact (that would be between 41.75 and 48.25%)
For the estimate
half-interval is z* sqrt(0.45*0.55/900)=0.0255
z^2*0.45*0.55/900=0.0255^2
z^2=2.36
z=+/-1.54
This would encompass 88% of the probability.